a car covers half of the distance with speed 60km/h and rest of distance wiyh 30km/h then the avg. speed will be
 

Grade:11

2 Answers

Shubham Singh
8 years ago
40 km/h
let t1 and t2 be time of journey with speed 60 and 30 respectively
total distance travelled = 60t1 + 30t2
and 60t1 = 30t2
total time = t1 + t2
now u can calculate avg. speed
noogler
8 years ago
avg speed=40kmph
 
use the formula Vavg=2v1v2/v1+v2
it is obtained by
 
a vehicle travels half of the distance with speed v1 in time t1and another half with speed v2 in time t2
v1=(L/2)/t1 , v2=(L/2)/t2.......................(1)
Vavg=total distance/total time=(L/2+L/2)/t1+t2=L/t1+t2
substitute t1 and t2 values from (1) in Vavg